OpenStudy (anonymous):

What is the slope of the line which passes through (2, 6) and (0, 2)?

Nnesha (nnesha):

do you know the slope formula

OpenStudy (anonymous):

y=mx+b? that one.. or the other one hahah

Nnesha (nnesha):

no this is linear equation the formula that we use to find slope

OpenStudy (sleepyhead314):

\[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } \]when given points (x1, y1) and (x2, y2)

OpenStudy (anonymous):

oh okay

Nnesha (nnesha):

put the order pair in to the formula and get the answer

Nnesha (nnesha):

did you get the answer if yes so let me know the answer please :)

OpenStudy (anonymous):

-2 as the answer although i do not think it is right

OpenStudy (sleepyhead314):

not quite :) try doing 6-2 ----- = ? 2-0

OpenStudy (anonymous):

so 4...

Nnesha (nnesha):

its not negative